OpenStudy (anonymous):

Solve the system by graphing. 3x+y-2=0 6x+2y=6

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

will give medal

OpenStudy (anonymous):

they are trying to trick you

OpenStudy (anonymous):

the question

OpenStudy (anonymous):

\[3x+y-2=0\]is the same as \[3x+y=2\] right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

now take \(3x+y=2\) and double everything you get \(6x+2y=4\) which is the same line

OpenStudy (anonymous):

now \(x\) and \(y\) are variables, they can be anything that fits the equation, but it is not possible for \(6x+2y\) to be both 4 and 6 at the same time

OpenStudy (anonymous):

infinte solution

OpenStudy (anonymous):

NO solution

OpenStudy (anonymous):

the two lines are parallel

OpenStudy (anonymous):

o so there not the same line

OpenStudy (anonymous):

no

OpenStudy (anonymous):

one is \(6x+2y=4\) and the other is \(6x+2y=6\)

OpenStudy (anonymous):

o